Chlorophyll the green colouring matter of plant contains 2.68% of magnesium. Calculate the number of atoms of Mg in 2g of Chlorophyll.

Grade:11

4 Answers

ABHISHEK JAIN AskiitiansExpert-IITD
13 years ago

Dear Chanchal,

since  chlorophyll contains 2.68% Mg therefore weight of Mg in 2g chlorophyll = 2×2.68/100 = .0536 g

now Mole of Mg = weight/atomic weight

                    = .0536/24

       since 1 mole contains  6.023×10^23 particles

        therefore number of atoms present in .0536/24 mole will be = (.0536/24) × 6.023×10^23

                                                                                               =1.34 × 10^23 atoms of Mg

 

 

We are all IITians and here to help you in your IIT JEE preparation.

 

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

YOUR NAME

Abhishek Jain

AskiitiansExpert-IITD

 

santosh kumar
13 years ago

the no. of mg atoms in 2g of chlorophyll =1300000000000000000000 atoms

Sachin
6 years ago
1.3×10^21 are total number of molecules in 2.00 g of C6H12O6(glucose).i am iitian iit mumbai Sachin Agarwalif any query ask me
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

since chlorophyll contains 2.68% Mg
therefore weight of Mg in 2g chlorophyll = 2×2.68/100
= .0536 g

now Mole of Mg = weight/atomic weight
= .0536/24
since 1 mole contains 6.023×10^23 particles
therefore number of atoms present in .0536/24 mole will be
= (.0536/24) × 6.023×10^23
=1.34 × 10^23 atoms of Mg

Thanks and Regards